# DualStreamFoveaNet: A Dual Stream Fusion Architecture with Anatomical Awareness for Robust Fovea Localization

### v4

@@ 4.2 Implementation Details and Evaluation Metrics
All experiments are implemented in PyTorch and conducted on one NVIDIA GeForce RTX TITAN GPU. We use the Adam optimizer [47] with a batch size of 2. The loss function for predicted segmentation map and ground-truth is a summation of dice loss and binary cross-entropy. The experimental setup for the DSFN architectures and the splitting of training and test sets on the Messidor and PALM datasets follow the identical approach as previously reported in [15, 14]. For our proposed DSFN, we set the initial learning rate to $1e^{-3}$, which is gradually decayed to $1e^{-9}$ using the CosineAnnealingLR strategy over 300 epochs on Messidor and PALM. For the Tisu dataset, we set the initial learning rate to $6e^{-5}$ for all related experiments.
5 Results
5.1 Comparison to State-of-the-Art
@@
${}^{a}$ The GPU usage when training (MiB)
${}^{b}$ The GPU usage when inferencing (MiB)
@@
In Table 2, we compare the performance of DSFN with existing methods on the public dataset, Messidor and PALM. Methods are classified based on whether they use deep learning techniques and whether they incorporate multi-cue features. We observe that most traditional morphological methods [16, 17, 2, 18, 19, 20] rely on landmarks outside the macula, such as vessels or the optic disc. However, in most deep learning-based studies, including coordinate regression methods [24, 26, 14] ($\checkmark/C$ in Table 2) and segmentation methods [25, 27] ($\checkmark/M$), only fundus images are used, resulting in poor incorporation of anatomical relationships throughout the entire image, leading to failure in more challenging cases. Bilateral-ViT [15] is the only previous deep learning-based method that incorporates fundus and vessel features for fovea localization, which achieves better accuracy and robustness. The proposed DSFN, which combines a transformer-based multi-cue fusion encoder and adaptive learning tokens, outperforms all previous studies in terms of fovea localization accuracy on the Messidor and PALM datasets. Specifically, in Table 2, DSFN achieves the highest accuracy of 98.86% at $1/4R$, with gains of 0.71% and 3.53% compared to previous works [14] and [27], respectively. Our network also achieves better performance than Bilateral-ViT [15] and its light version. At evaluation thresholds of $1/2R$, $1R$, and $2R$, DSFN achieves 100% accuracy, indicating a localization error of at most $1/2R$ (approximately 19 pixels for an input image size of $512\times 512$). The PALM dataset is more challenging, with a smaller number of images and complex diseased patterns. Our proposed DSFN demonstrates superiority over all other methods on this dataset in Table 2, achieving accuracies of 69% and 85% at $1/4R$ and $1/2R$, respectively, which are 4% and 2% better than Bilateral-ViT. In addition, our method achieves a 14% improvement ($1/4R$) over Bilateral-ViT/Lit, and a 3% improvement ($1R$) over both Bilateral-ViT/Lit and [14]. Therefore, our DSFN achieves state-of-the-art performance on both Messidor and PALM datasets with high computational efficiency.
@@ 5.2 Fovea Localization on Normal and Diseased Images
In Table 3, we separately evaluate the performance of fovea localization for normal and diseased cases in Messidor and PALM datasets to assess the robustness of our method. We compare our proposed DSFN to several widely used segmentation networks, including UNet [31], U2 Net [34], and a hybrid version of TransUNet with ResNet50 for patch embedding [42]. The models of Bilateral-ViT, Bilateral-ViT/Lit and our DSFN are identical to those used in Table 2. In Table 3, the proposed DSFN achieves the lowest error (Err) on both datasets. Bilateral-ViT [15] and DSFN both obtain 100% accuracy from $1/2R$ to $2R$ on all the Messidor images, and achieve 100% accuracy of $1R$ and $2R$ on normal PALM images. Compared to existing methods, DSFN demonstrates superior performance on almost all metrics on Messidor. As shown in Table 1 and Table 3, although DSFN has only 25% FLOPs (62.11G) compared to Bilateral-ViT (249.89G), it achieves better performance on diseased images of PALM from $1/4R$ to $2R$, with up to 7.54% improvement compared to the other methods ($1/4R$, Diseased). Its improvement is significantly increased to 18.86% ($1/4R$, Diseased) compared to Bilateral-ViT/Lit (83.05G, i.e., the network with the closest FLOPs to DSFN). Thus, our proposed DSFN is highly reliable in fovea localization on both normal and diseased fundus images with high computational efficacy.
@@ 5.3 Comparison of Multi-Cue Fusion Architectures
To comprehensively assess the performance of models with input features from multiple cues (fundus and vessel distributions), we implement a multi-cue fusion version for the baseline models, UNet, U2 Net and TransUNet. We utilize two identical encoders, each with an input of a fundus image and a vessel map. The features are extracted independently and concatenated at the bottleneck for decoding (similar to [48, 28]). These modified baseline models are referred to as UNet-MF, U2 Net-MF and TransUNet-MF. The results in Fig. 3 show that architectures using multi-cue features outperform typical networks with fundus-only input. This is evident in the challenging datasets, PALM and Tisu, where the improvement is pronounced. Moreover, the Tisu dataset is more complex than PALM, with more images (4103 vs. 400) and a wider range of disease types and severity. Therefore, the results on PALM and Tisu demonstrate the potential of architectures that can effectively handle complex datasets. Fig. 3 shows a comparison of the described architectures (mean error against FLOPs) on PLAM (red markers) and Tisu (blue markers). Below each marker, we provide the corresponding mean error, and dashed lines connect each standard baseline model with its multi-cue fusion architecture (MF). The multi-cue fusion versions (UNet-MF, U2 Net-MF, and TransUNet-MF) outperform their standard versions at a considerably higher computational cost due to the additional encoder. For all multi-cue fusion architectures, the proposed DSFN achieves the best results with the smallest errors on both PLAM (48.72 pixels) and Tisu (36.46 pixels). Furthermore, it demands only 62.11G FLOPs, which is four times less than the FLOPs required by the next best-performing approach, Bilateral-ViT (249.89G). Compared to the model with comparable FLOPs (Bilateral-ViT/Lit, 83.05G), DSFN shows significant advantages of 13.75 and 3.92 on PLAM and Tisu, respectively. Fig. 4 provides visual results of fovea localization on images with severe diseases from the PALM and Tisu datasets. These images in Fig. 4-a and Fig. 4-b suffer from poor image quality and non-standard fovea locations, respectively. Our DSFN generates the most accurate predictions for several challenging cases with poor lighting conditions and blurred appearance (Fig. 4-a). For another challenging types (Fig. 4-b), where the macula is close to the image boundary, the predictions of DSFN (green crosses) are closest to the ground-truth (white crosses). On the contrary, predictions from alternative architectures lacking the ability to globally incorporate long-range multi-cue features may lead to fovea locations appearing on the incorrect side of the optic disc (Fig. 4-b). These results suggest that DSFN can adequately model fundus and vessel features of two streams, resulting in superior performance compared to other networks.
@@ 5.5.1 Comparison of Inputs for DSFN
Table 5 compares the performance of DSFN when using different inputs. The standard input configuration uses fundus images and vessel maps as inputs for the main and satellite streams, respectively (Fundus+Vessel). These experiments achieve the best accuracy on all metrics, with the smallest mean error on both PALM (48.72 pixels) and Tisu (36.46 pixels). When using fundus images as the second input (Fundus+Fundus), the model’s performance slightly degrades as feeding fundus images into the satellite stream does not provide the explicit anatomical structure for the BTI module to learn where to focus its attention. Experiments using vessel maps as both inputs (Vessel+Vessel) lead to significant accuracy decreases on both PALM and Tisu by 23.53 and 11.67 pixels, respectively, indicating a severe loss of information. In addition, we conduct experiments where the input of the satellite stream (i.e., vessel distribution map) is pre-trained and generated using a weaker architecture, UNet ($*{}^{*}$), as opposed to TransUNet in the default setting (Fundus+Vessel). These experiments on PALM and Tisu demonstrate the performance of our DSFN when the vessel input serves as a weak cue. The outcomes presented in Table 5 suggest that the performance of $*{}^{*}$ generally falls between that of Fundus+Vessel and Fundus-only. This finding demonstrates that even when providing a weak cue, the DSFN effectively learns valuable anatomical-related features, instead of starting to learn spatial attention from scratch (i.e., Fundus-only).
PALM ${}_{\downarrow}$ 1/4 R (%) 1/2 R (%) 1R (%) 2R (%)
Fundus+Vessel (Ours) 48.72 69 85 97 98
Fundus-only 54.46 64 85 95 98
Vessel-only 72.25 57 75 92 97
$*{}^{*}$ 49.61 67 84 95 98
Tisu ${}_{\downarrow}$ 1/4 R (%) 1/2 R (%) 1R (%) 2R (%)
Fundus+Vessel (Ours) 36.46 52.32 75.49 92.93 97.44
Fundus-only 37.48 52.32 73.78 91.10 96.95
Vessel-only 48.13 33.66 61.83 89.39 96.83
$*{}^{*}$ 37.34 51.59 73.66 91.95 97.44
*** denotes that the input of the satellite stream is pre-trained and
generated by UNet (i.e., a weaker cue).
We utilize spatial attention maps to elucidate the self-learning features extracted by TokenLearners, visually demonstrating these weight maps in Fig. 5. To visualize the attention of tokens in an element-wise manner, we maximize the saliency values among all the tokens $n$ for each position ${h\times w}$ of the spatial attention map $\text{F}_{\text{attn}}\in\mathbb{R}^{h\times w\times n}$ (refer to Eq. 2). Subsequently, we normalize these values and superimpose the resulting maps onto the corresponding fundus, facilitating a comparison of their structural relationships. In our experiments using Fundus+Vessel inputs, the spatial attention maps focus on structural features along the optic disc and the direction of vessel branches (Fig. 5-a). This focus is feasible due to the significant anatomical relationships between these structures and the foveal region[9, 2, 10, 7, 8]. In contrast, despite having more detailed information in Fundus+Fundus experiments, TokenLearners, which are not guided by explicit anatomical structures, fail to learn features along the vessels (Fig. 5-b). This results in fewer tokens carrying useful features for fovea localization and may limit the effectiveness of the intermediate BTI modules in DSFN, leading to a slight underperformance on PALM and Tisu datasets (Table 5). Therefore, the structural information provided by vessel maps as the second input is crucial for achieving accurate fovea localization.
@@
Figure 5: Visualizations generated from spatial attention maps, indicating the focus of TokenLearner in BTI module. These visual results have been resized and superimposed onto the the corresponding fundus. Circles with dotted lines represent the position of optic discs, and arrows represents the direction of main vessel branches.
@@ 5.5.2 Comparison of Methods for Reducing and Recovering Tokens
To evaluate the effectiveness of different components within our BTI module for reducing and recovering token numbers, we perform a comprehensive set of ablation experiments on the PALM and Tisu datasets. Instead of employing TokenLearner (TL) and TokenFuser (TF) with adaptively learnable parameters, we alternatively test more straightforward methods used in [28], average pooling (AvgPool) and bilinear interpolation (Interpolate), respectively. In Fig. 6, we demonstrate a visualization of the mean error against the computational cost (FLOPs) on PALM (red) and Tisu (blue) datasets. Experiments using both TokenLearner and TokenFuser (TL+TF) achieve the best performance on PALM and Tisu with mean errors of 48.72 and 36.46 pixels, respectively. This is in contrast to using more straightforward methods such as average pooling (AvgPool) and bilinear interpolation (Interpolate), which may lead to a loss of information and reduced performance. In the proposed DSFN (TL+TF), total excess costs of FLOPs for the four BTI modules utilizing TokenLearner and TokenFuser are only 0.85G and 0.61G, respectively. This slight increase in computation leverages significant performance benefits, demonstrating the high efficacy of the adaptively learnable parameters of TokenLearner and TokenFuser in our architecture.

### v5

@@ 4.2 Implementation Details and Evaluation Metrics
a The GPU usage when training (MiB)
b The GPU usage when inferencing (MiB)
All experiments are implemented in PyTorch and conducted on one NVIDIA GeForce RTX TITAN GPU. We use the Adam optimizer [56] with a batch size of 2. The loss function for predicted segmentation map and ground-truth is a summation of dice loss and binary cross-entropy. The experimental setup for the DSFN architectures and the splitting of training and test sets on the Messidor and PALM datasets follow the identical approach as previously reported in [15, 14]. For our proposed DSFN, we set the initial learning rate to $1e^{-3}$, which is gradually decayed to $1e^{-9}$ using the CosineAnnealingLR strategy over 300 epochs on Messidor and PALM. For the Tisu dataset, we set the initial learning rate to $6e^{-5}$ for all related experiments. To reduce overfitting, we adopt a data augmentation strategy to increase the diversity of training data (Details can be found in APPENDIX B). In addition, Xavier weight initialization is utilized for MLP weights in MHSA layers, and normal weight initialization is utilized for MLP biases in MHSA layers.
@@
5 Results
5.1 Comparison to State-of-the-Art
In Table 2, we compare the performance of DSFN with existing methods on the public dataset, Messidor and PALM. Methods are classified based on whether they use deep learning techniques and whether they incorporate multi-cue features. We observe that most traditional morphological methods [16, 17, 2, 18, 19, 20] rely on landmarks outside the macula, such as vessels or the optic disc. However, in most deep learning-based studies, including coordinate regression methods [25, 27, 14] ($\checkmark/C$ in Table 2) and segmentation methods [26, 28] ($\checkmark/M$), only fundus images are used, resulting in poor incorporation of anatomical relationships throughout the entire image, leading to failure in more challenging cases. Bilateral-ViT [15] is the only previous deep learning-based method that incorporates fundus and vessel features for fovea localization, which achieves better accuracy and robustness.
The proposed DSFN, which combines a transformer-based multi-cue fusion encoder and adaptive learning tokens, outperforms all previous studies in terms of fovea localization accuracy on the Messidor and PALM datasets. Specifically, in Table 2, DSFN achieves the highest accuracy of 98.86% at $1/4R$, with gains of 0.71% and 3.53% compared to previous works [14] and [28], respectively. Our network also achieves better performance than Bilateral-ViT [15] and its light version. At evaluation thresholds of $1/2R$, $1R$, and $2R$, DSFN achieves 100% accuracy, indicating a localization error of at most $1/2R$ (approximately 19 pixels for an input image size of $512\times 512$).
The PALM dataset is more challenging, with a smaller number of images and complex diseased patterns. Our proposed DSFN demonstrates superiority over all other methods on this dataset in Table 2, achieving accuracies of 69% and 85% at $1/4R$ and $1/2R$, respectively, which are 4% and 2% better than Bilateral-ViT. In addition, our method achieves a 14% improvement ($1/4R$) over Bilateral-ViT/Lit, and a 3% improvement ($1R$) over both Bilateral-ViT/Lit and [14]. Therefore, our DSFN achieves state-of-the-art performance on both Messidor and PALM datasets with high computational efficiency.
@@ 5.2 Fovea Localization on Normal and Diseased Images
In Table 3, we separately evaluate the performance of fovea localization for normal and diseased cases in Messidor and PALM datasets to assess the robustness of our method. We compare our proposed DSFN to several widely used segmentation networks, including UNet [37], U2 Net [41], and a hybrid version of TransUNet with ResNet50 for patch embedding [49]. The models of Bilateral-ViT, Bilateral-ViT/Lit and our DSFN are identical to those used in Table 2.
In Table 3, the proposed DSFN achieves the lowest error (Err) on both datasets. Bilateral-ViT [15] and DSFN both obtain 100% accuracy from $1/2R$ to $2R$ on all the Messidor images, and achieve 100% accuracy of $1R$ and $2R$ on normal PALM images. Compared to existing methods, DSFN demonstrates superior performance on almost all metrics on Messidor. As shown in Table 1 and Table 3, although DSFN has only 25% FLOPs (62.11G) compared to Bilateral-ViT (249.89G), it achieves better performance on diseased images of PALM from $1/4R$ to $2R$, with up to 7.54% improvement compared to the other methods ($1/4R$, Diseased). Its improvement is significantly increased to 18.86% ($1/4R$, Diseased) compared to Bilateral-ViT/Lit (83.05G, i.e., the network with the closest FLOPs to DSFN). Thus, our proposed DSFN is highly reliable in fovea localization on both normal and diseased fundus images with high computational efficacy.
@@ 5.3 Comparison of Multi-Cue Fusion Architectures
To comprehensively assess the performance of models with input features from multiple cues (fundus and vessel distributions), we implement a multi-cue fusion version for the baseline models, UNet, U2 Net and TransUNet. We utilize two identical encoders, each with an input of a fundus image and a vessel map. The features are extracted independently and concatenated at the bottleneck for decoding (similar to [58, 29]). These modified baseline models are referred to as UNet-MF, U2 Net-MF and TransUNet-MF. The results in Fig. 4 show that architectures using multi-cue features outperform typical networks with fundus-only input. This is evident in the challenging datasets, PALM and Tisu, where the improvement is pronounced. Moreover, the Tisu dataset is more complex than PALM, with more images (4103 vs. 400) and a wider range of disease types and severity. Therefore, the results on PALM and Tisu demonstrate the potential of architectures that can effectively handle complex datasets.
@@
Fig. 4 shows a comparison of the described architectures (mean error against FLOPs) on PLAM (red markers) and Tisu (blue markers). Below each marker, we provide the corresponding mean error, and dashed lines connect each standard baseline model with its multi-cue fusion architecture (MF). The multi-cue fusion versions (UNet-MF, U2 Net-MF, and TransUNet-MF) outperform their standard versions at a considerably higher computational cost due to the additional encoder. For all multi-cue fusion architectures, the proposed DSFN achieves the best results with the smallest errors on both PLAM (48.72 pixels) and Tisu (36.46 pixels). Furthermore, it demands only 62.11G FLOPs, which is four times less than the FLOPs required by the next best-performing approach, Bilateral-ViT (249.89G). Compared to the model with comparable FLOPs (Bilateral-ViT/Lit, 83.05G), DSFN shows significant advantages of 13.75 and 3.92 on PLAM and Tisu, respectively.
Fig. 5 provides visual results of fovea localization on images with severe diseases from the PALM and Tisu datasets. These images in Fig. 5-a and Fig. 5-b suffer from poor image quality and non-standard fovea locations, respectively. Our DSFN generates the most accurate predictions for several challenging cases with poor lighting conditions and blurred appearance (Fig. 5-a). For another challenging types (Fig. 5-b), where the macula is close to the image boundary, the predictions of DSFN (green crosses) are closest to the ground-truth (white crosses). On the contrary, predictions from alternative architectures lacking the ability to globally incorporate long-range multi-cue features may lead to fovea locations appearing on the incorrect side of the optic disc (Fig. 5-b). These results suggest that DSFN can adequately model fundus and vessel features of two streams, resulting in superior performance compared to other networks.
PALM Err 1/4 R 1/2 R 1R 2R (%)
Fundus+Vessel (Ours) 48.72 69 85 97 98
Fundus-only 54.46 64 85 95 98
Vessel-only 72.25 57 75 92 97
Fundus+Vessel 49.61 67 84 95 98
Tisu Err 1/4 R 1/2 R 1R 2R (%)
Fundus+Vessel (Ours) 36.46 52.32 75.49 92.93 97.44
Fundus-only 37.48 52.32 73.78 91.10 96.95
Vessel-only 48.13 33.66 61.83 89.39 96.83
Fundus+Vessel 37.34 51.59 73.66 91.95 97.44
* denotes that the input of the satellite stream is pre-trained and
generated by UNet (i.e., a weaker cue).
@@ 5.5.1 Comparison of Inputs for DSFN
Table 5 compares the performance of DSFN when using different inputs. The standard input configuration uses fundus images and vessel maps as inputs for the main and satellite streams, respectively (Fundus+Vessel). These experiments achieve the best accuracy on all metrics, with the smallest mean error on both PALM (48.72 pixels) and Tisu (36.46 pixels). When using fundus images as the second input (Fundus+Fundus), the model’s performance slightly degrades as feeding fundus images into the satellite stream does not provide the explicit anatomical structure for the BTI module to learn where to focus its attention. Experiments using vessel maps as both inputs (Vessel+Vessel) lead to significant accuracy decreases on both PALM and Tisu by 23.53 and 11.67 pixels, respectively, indicating a severe loss of information. In addition, we conduct experiments where the input of the satellite stream (i.e., vessel distribution map) is pre-trained and generated using a weaker architecture, UNet (Fundus+Vessel), as opposed to TransUNet in the default setting (Fundus+Vessel). These experiments on PALM and Tisu demonstrate the performance of our DSFN when the vessel input serves as a weak cue. The outcomes presented in Table 5 suggest that the performance of Fundus+Vessel generally falls between that of Fundus+Vessel and Fundus-only. This finding demonstrates that even when providing a weak cue, the DSFN effectively learns valuable anatomical-related features, instead of starting to learn spatial attention from scratch (i.e., Fundus-only).
@@
Figure 6: Visualizations generated from spatial attention maps, indicating the focus of TokenLearner in BTI module. These visual results have been resized and superimposed onto the the corresponding fundus. Circles with dotted lines represent the position of optic discs, and arrows represents the direction of main vessel branches. The white, green, and red crosses in the first column are the ground-truth, prediction result with the second input of vessel (a) and fundus (b), respectively.
We utilize spatial attention maps to elucidate the self-learning features extracted by TokenLearners, visually demonstrating these weight maps in Fig. 6. To visualize the attention of tokens in an element-wise manner, we maximize the saliency values among all the tokens $n$ for each position ${h\times w}$ of the spatial attention map $\text{F}_{\text{attn}}\in\mathbb{R}^{h\times w\times n}$ (refer to Eq. 2). Subsequently, we normalize these values and superimpose the resulting maps onto the corresponding fundus, facilitating a comparison of their structural relationships.
In our experiments using Fundus+Vessel inputs, the spatial attention maps focus on structural features along the optic disc and the direction of vessel branches (Fig. 6-a). This focus is feasible and consistent to the significant anatomical relationships between these structures and the fovea region (shown in Fig. 1[9, 2, 10, 8, 7]. In contrast, despite having more detailed information in Fundus+Fundus experiments, TokenLearners, which are not guided by explicit anatomical structures, fail to learn features along the vessels (Fig. 6-b). This results in fewer tokens carrying useful features for fovea localization and may limit the effectiveness of the intermediate BTI modules in DSFN, leading to a slight underperformance on PALM and Tisu datasets (Table 5). In the first column of Fig. 6, the predictions of Fundus+Vessel (green crosses) show superior results compared to those of Fundus+Fundus (red crosses). Therefore, the structural information provided by vessel maps as the second input is crucial for achieving accurate fovea localization.
@@ 5.5.2 Comparison of Methods for Reducing and Recovering Tokens
Table 6: Comparison of performance between different configurations of learned tokens and layer number of MHSA blocks utilized in the BTI module. The best results are highlighted in bold.
Optimal Tokens Layers Err 1/4 R 1/2 R 1R 2R (%)
8 12 54.40 60 81 93 97
$4\times 4$ 12 52.45 61 78 93 97
$8\times 8$ 12 48.72 69 85 97 98
$16\times 16$ 12 52.66 64 84 94 98
$8\times 8$ 8 49.96 66 82 96 98
$8\times 8$ 16 55.06 68 85 95 97
To evaluate the effectiveness of different components within our BTI module for reducing and recovering token numbers, we perform a comprehensive set of ablation experiments on the PALM and Tisu datasets. Instead of employing TokenLearner (TL) and TokenFuser (TF) with adaptively learnable parameters, we alternatively test more straightforward methods used in [29], average pooling (AvgPool) and bilinear interpolation (Interpolate), respectively.
In Fig. 7, we demonstrate a visualization of the mean error against the computational cost (FLOPs) on PALM (red) and Tisu (blue) datasets. Experiments using both TokenLearner and TokenFuser (TL+TF) achieve the best performance on PALM and Tisu with mean errors of 48.72 and 36.46 pixels, respectively. This is in contrast to using more straightforward methods such as average pooling (AvgPool) and bilinear interpolation (Interpolate), which may lead to a loss of information and reduced performance. In the proposed DSFN (TL+TF), total excess costs of FLOPs for the four BTI modules utilizing TokenLearner and TokenFuser are only 0.85G and 0.61G, respectively. This slight increase in computation leverages significant performance benefits, demonstrating the high efficacy of the adaptively learnable parameters of TokenLearner and TokenFuser in our architecture.
5.5.3 Comparison of Hyperparameter Settings
The above subsections have evaluated the necessity of vessel distribution inputs and of the BTI module. We also conduct ablation studies to evaluate, in the BTI module, how the performance of feature integration and fusion is affected by the structure of the MHSA blocks, to ensure optimal configurations.
Table 6 presents a comparative analysis of the performance on PALM between different configurations of the number of learned tokens and Multi-Head Self-Attention (MHSA) layers utilized in the BTI module. The table evaluates various configurations based on several metrics: mean error, and success rates at different thresholds (1/4 R, 1/2 R, 1R, and 2R). The optimal configuration identified is the use of $8\times 8$ learned tokens with 12 MHSA layers, which achieves the best results across all metrics, including the lowest mean error (48.72) and the highest success rates at all thresholds. Other configurations, such as $4\times 4$ tokens with 12 layers and $8\times 8$ tokens with 8 layers, show lower performance, indicating that both the number of tokens and the depth of MHSA layers significantly influence the performance of feature integration and the accuracy of fovea localization.